### 07 Apr 2026 v2

# Neural-NPV Control: Learning Parameter-Dependent Controllers and Lyapunov Functions with Neural Networks

Abstract
Nonlinear parameter-varying (NPV) systems are a class of nonlinear systems whose dynamics explicitly depend on time-varying external parameters, making them suitable for modeling real-world systems with dynamics variations. Traditional synthesis methods for NPV systems, such as sum-of-squares (SOS) optimization, are only applicable to control-affine systems, face scalability challenges and often lead to conservative results due to structural restrictions. To address these limitations, we propose Neural-NPV, a two-stage learning-based framework that leverages neural networks to jointly synthesize a PD controller and a PD Lyapunov function for an NPV system under input constraints. In the first stage, we utilize a computationally cheap, gradient-based counterexample-guided procedure to synthesize an approximately valid PD Lyapunov function and a PD controller. In the second stage, a level-set guided refinement is then conducted to obtain a valid Lyapunov function and controller while maximizing the robust region of attraction (R-ROA). We demonstrate the advantages of Neural-NPV in terms of applicability, performance, and scalability compared to SOS-based methods through numerical experiments involving an simple inverted pendulum with one scheduling parameter and a quadrotor system with three scheduling parameters.
I Introduction
A nonlinear parameter-varying (NPV) system is a nonlinear system whose behavior explicitly depends on external time-varying parameters. This concept extends the linear parameter-varying (LPV) framework [17], by allowing for the dynamics to have a nonlinear instead of a linear structure. Many real-world nonlinear time-varying systems can be effectively represented as NPV systems. Examples include rockets with decreasing mass due to fuel consumption, ground vehicles subjected to variable friction coefficients, and aircraft experiencing disturbances that change over time. Similar to the LPV framework [17, 12], representing a nonlinear time-varying system as an NPV system facilitates the design of gain-scheduled nonlinear controllers that adjust the control strategy dynamically based on real-time measurements and estimates of the scheduling parameters, while the limitations of the LPV framework resulting from local validity or over-approximation of the LPV models [17].
Recent studies have explored the synthesis of nonlinear parameter-dependent (PD) controllers for NPV systems using SOS programming [8, 10, 26]. However, SOS programming is not scalable to high-dimensional systems. Additionally, SOS-based nonlinear synthesis often imposes a restrictive structure on the Lyapunov functions to avoid a non-convex problem, leading to potentially conservative results. To overcome the limitations of SOS-based NPV synthesis, this paper investigates the use of neural networks to jointly learn PD controllers and Lyapunov functions for NPV systems.
I-A Related work
LPV control: In the LPV control approach [19, 17], a nonlinear system is approximated by a LPV model. Using this LPV representation, controllers can be developed to achieve stability and performance guarantees for the closed-loop system, often through the use of linear matrix inequalities (LMIs) [17]. However, a key limitation of this methodology is that the LPV model may only be accurate locally when it is derived via Jacobian linearization. Alternatively, when the quasi-LPV modeling technique is applied, the model tends to be a conservative over-approximation of the actual nonlinear dynamics, which can restrict controller performance [17].
SOS and its use in NPV synthesis: One of the most widely used methods for nonlinear analysis and synthesis is SOS optimization [13]. It has been applied to jointly synthesize nonlinear parameter-dependent (PD) controllers and PD Lyapunov functions for NPV systems [8, 10, 26]. However, SOS optimization-based methods suffer from two major limitations: (i) Restricted scalability: The computational complexity grows rapidly with the dimensionality of the system and the degree of the polynomial basis used, rendering them impractical for high-dimensional systems [2]. (ii) Conservative performance due to structural restrictions: SOS-based nonlinear synthesis typically imposes structures on certificate functions to avoid a nonconvex problem, leading to conservative performance. For example, SOS-based Lyapunov synthesis typically requires that the Lyapunov functions do not depend on states whose derivatives are directly affected by control inputs [14, 26].
Figure 1: Neural-NPV framework. Within Neural-NPV, a PD controller $\pi(x,\theta)$ (to stabilize the NPV system) and a PD Lyapunov function $V(x,\theta)$ (to certify the stability of the closed-loop system) are jointly trained to minimize the Lyapunov loss via supervised learning. ROI: Region of interest; ROA: Region of attraction
Learning certified control: There has been an increasing trend for learning a control certificate (alongside a controller) using NNs (NNs) [5]. Such a neural certificate can be a Lyapunov function for certifying the stability of an equilibrium point [15, 3], or a contraction metric for certifying stability about trajectories [20, 21] or a barrier function for certifying set invariance [16, 7], among others. Compared to SOS programming-based nonlinear synthesis, certified neural control can be applied to general nonlinear systems (in contrast to polynomial systems for SOS programming-based methods), and is scalable to high-dimensional systems. It also allows for exploring more general forms of controllers and certificates, potentially reducing the conservatism of traditional methods that typically impose special structures on the controllers and/or certificates (e.g., Lyapunov functions) to avoid a non-convex problem [14, 11, 24, 18, 26]. For a comprehensive overview of these methods, readers can refer to the survey paper [5]. Compared to traditional optimization-based methods for control certificate synthesis, such as sum of squares (SOS) programming [13, 14], Nevertheless, existing work typically consider time-invariant dynamics [6]. To the best of our knowledge, there is no study on learning-based control for NPV systems.
I-B Our contribution
The key contributions of our work are as follows:
  • We propose Neural-NPV, a novel two-stage framework for jointly synthesizing a PD Lyapunov function and a PD controller using neural networks for general NPV systems, unlike existing SOS-based methods that require a control-affine structure or polynomial parameterizations for the dynamics. Our framework can also handle asymmetric input constraints, which is challenging for SOS-based methods [8, 25]. We verify the learned PD Lyapunov function and PD controller through two empirical verification schemes.
  • We show the superior performance of our framework in yielding a much larger robust region of attraction (R-ROA) compared to SOS-based methods using numerical experiments. We also demonstrate the scalability of our approach through numerical experiments on a quadrotor system with three scheduling parameters denoting the disturbances.
II Problem settings and preliminaries
We consider a nonlinear parameter-varying (NPV) system,
$\dot{x}=f(x,u,\theta),$ (1)
where $x(t)\in\mathbb{R}^{n}$ is the state vector, $u(t)\in\mathbb{R}^{n_{u}}$ is input vector, $\theta(t)\in\mathbb{R}^{n_{\theta}}$ denotes the time-varying parameter vector that can be measured online and $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\times\mathbb{R}^{n_{u}}\times\mathbb{R}^{n_{\theta}}\ \rightarrow\mathbb{R}^{n}$ is locally Lipschitz in $x,u$ and $\theta$. The parameter $\theta(t)$ characterizes the variation of the dynamics. For instance, it can represent the mass of a rocket, which will be time-varying due to propellant consumption, or describe the external disturbance applied to an aircraft or a drone.
The control input is bounded as
$\underaccent{\bar}{u}\leq u(t)\leq\bar{u},\quad\forall t\geq 0,$ (2)
where $\underaccent{\bar}{u}\in\mathbb{R}^{m}$ and $\bar{u}\in\mathbb{R}^{m}$ denote the lower and upper bounds on $u$, respectively.
Assumption 1.
The time-varying parameters $\theta$ and their derivatives $\dot{\theta}$ satisfy
$\theta(t)\in\Theta\;\;\text{and}\;\;\dot{\theta}(t)\in{\Omega},\quad\forall\,t\geq 0,$ (3)
where $\Theta$ is a compact set and $\Omega$ is a hyper-rectangular set defined by,
$\Omega\triangleq\bigl\{\,v\in\mathbb{R}^{n_{\theta}}\;\mathrel{\mathop{\ordinarycolon}}\;\underline{v}_{i}\leq v_{i}\leq\overline{v}_{i},\;i\in\mathbb{Z}_{1}^{n_{\theta}}\,\bigr\}.$ (4)
For clarity, we define $\mathcal{F}_{\Theta}^{\Omega}$ as the set of all feasible parameter trajectories $\theta(t)$,
$\mathcal{F}^{\Omega}_{\Theta}\triangleq\Bigl\{\theta\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{+}\to\mathbb{R}^{n_{\theta}}\;\big|\theta(t)\in\Theta,\dot{\theta}(t)\in\Omega,\;\forall\,t\geq 0\Bigr\}.$ (5)
Given 1 subject to Assumption 1, we are interested in designing a nonlinear PD controller
$u=\pi(x,\theta),$ (6)
that will stabilize the system 1 under all admissible trajectories of $\theta$ and a PD Lyapunov function $V(x,\theta)$ certifying the stability. Under the controller 6, the closed-loop system is given by
$\dot{x}=f(x,\pi(x,\theta),\theta)\triangleq\bar{f}(x,\theta).$ (7)
Without loss of generality, we assume the $x^{*}=0$ is an equilibrium state of the closed-loop system.
Definition 1.
[9, Definition 4.5] The equilibrium point $x^{*}=0$ of the closed-loop system 7 is exponentially stable if there exist positive constants c,kc,~k, and $\lambda$ such that
$\hskip-8.53581pt\mathinner{\!\left\lVert x(t)\right\rVert}\!\leq\!k\mathinner{\!\left\lVert x(t_{0})\right\rVert}e^{-\lambda(t-t_{0})},\ \forall t\geq t_{0}\!\geq 0,\ \forall\mathinner{\!\left\lVert x(t_{0})\right\rVert}\!<\!c,$ (8)
for all admissible trajectories of $\theta$ satisfying 3.
According to Lyapunov stability theory (as demonstrated in the proof of Theorem 1 in [26]), if there exists a PD controller $u=\pi(x,\theta)$ and a PD continuously differentiable function $V(x,\theta)\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}$ such that for all $x\in\mathcal{X}$ with $\mathcal{X}$ being a region of interest (ROI),
$V(0,\theta)=0,\quad\forall\theta\in\Theta,$ (9a)
$k_{1}\|x\|^{2}\leq V(x,\theta)\leq k_{2}\|x\|^{2},\quad\forall\theta\in\Theta,$ (9b)
$\dot{V}(x,{\theta},\dot{\theta})\leq-k_{3}\|x\|^{2},\quad\forall\theta\in\Theta,\;\forall\dot{\theta}\in\textup{Ver}(\Omega),$ (9c)
where
$\dot{V}(x,{\theta},\dot{\theta})\!=\!\frac{\partial V}{\partial x}\cdot\bar{f}+\frac{\partial V}{\partial\theta}\cdot\dot{\theta},$ (10)
$\textup{Ver}(\Omega)$ denotes the set of vertices of $\Omega$, and $k_{1},~k_{2}$, and $k_{3}$ are positive constants, then the closed-loop system 7 is exponentially stable.
Definition 2.
A robust region of attraction (R-ROA) for the closed-loop system (6) under the controller $u=\pi(x,\theta)$ is the set of initial states $x(0)$ from which the closed-loop trajectories satisfy $\lim_{t\to\infty}x(t)=0$ for all admissible trajectories of $\theta$. Formally, an R-ROA is a set defined by,
$\mathcal{A}=\left\{x(0)\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}\ \forall\theta\in\mathcal{F}^{\Omega}_{\Theta},\ \lim_{t\to\infty}x(t)=0\right\}.$ (11)
Remark 1.
Since 9c is affine with respect to $\dot{\theta}$, if it holds for all $\theta\in\textup{Ver}(\Omega)$, then it holds for $\theta\in\Omega$.
III Neural-NPV: Learning PD Controllers and Lyapunov functions
We now present a method to jointly learn the PD controller and PD Lyapunov functions using NNs that (approximately) satisfy the conditions 9.
III-A Parameterizing the PD controller and Lyapunov function
The Lyapunov function should meet all the conditions in 9. To simplify the training process, we parameterize our Lyapunov function in a way that conditions 9a & 9b can be automatically satisfied. Inspired by [23], we adopt the following parameterization:
$V(x,\theta)=x^{T}\left(\phi_{\textup{NN}}(x,\theta)\phi_{\textup{NN}}^{T}(x,\theta)+\epsilon I\right)x,$ (12)
where $\epsilon$ is a small positive constant, $I$ is an identity matrix, and $\phi_{\textup{NN}}(x,\theta)$ is an NN with $n+n_{\theta}$ neurons in the input layer and $n$ neurons in the output layer. $\phi_{\textup{NN}}(x,\theta)$ takes both state $x$ and parameters $\theta$ as the input. We use tanh activation function for $\phi_{\textup{NN}}$, which allows the computation of the derivative of $V(x,\theta)$ defined in 10.
With the parameterization in 12, 9a can be trivially satisfied. Also, $\phi_{\textup{NN}}(x,\theta)\phi_{\textup{NN}}^{T}(x,\theta)$ is positive semi-definite. Letting $\bar{\lambda}\geq 0$ and $\underaccent{\bar}{\lambda}\geq 0$ denote the largest and smallest eigenvalues of $\phi_{\textup{NN}}(x,\theta)\phi_{\textup{NN}}^{T}(x,\theta)$ for all $x\in\mathcal{X}$ and $\theta\in\Theta$, respectively. Then, we have
$(\underaccent{\bar}{\lambda}+\epsilon)\mathinner{\!\left\lVert x\right\rVert}^{2}\leq V(x,\theta)\leq(\bar{\lambda}+\epsilon)\mathinner{\!\left\lVert x\right\rVert}^{2},\ \forall x\in\mathcal{X},\theta\in\Theta.$ (13)
Setting $k_{1}=\underaccent{\bar}{\lambda}+\epsilon>0$ and $k_{2}=\bar{\lambda}+\epsilon>0$, 13 implies that 9b is satisfied for all $x\in\mathcal{X}$.
To represent the controller, we use a multi-layer NN $\pi_{\textup{NN}}(x,\theta)$ with $n+n_{\theta}$ neurons in the input layer and $m$ neurons in the output layer. We denote the weight parameters of $\phi_{\textup{NN}}(x,\theta)$ and $\pi_{\textup{NN}}(x,\theta)$ as $W_{V}$ and $W_{\pi}$, respectively. To ensure the controller generates desired input $u(x^{*},\theta)$ at the equilibrium state $x^{*}$ which achieves $\dot{x}^{*}=\bar{f}(x^{*},\theta)=0$ for all $\theta\in\Theta$, we parameterize the normalized controller as
$\tilde{u}(x,\theta)=\pi_{\textup{NN}}(x,\theta)-\pi_{\textup{NN}}(x^{*},\theta)+\tilde{u}(x^{*},\theta).$ (14)
. To enforce the constraint 2, we use tanh in the output layer of the controller $\pi_{\textup{NN}}$ to restrict the normalized output within [1,1][-1,1]. $\tilde{u}(x,\theta)$ and $\tilde{u}(x^{*},\theta)$ are normalized value of $u(x,\theta)$ and $u(x^{*},\theta)$, respectively. Rather than using projection or clipping, which introduces complications in gradient-based training, we employ a smooth, differentiable activation at the output, which empirically improves the training process. We scale the normalized output to actual physical control input using an affine de-normalization to get the final controller input:
$u=\pi(x,\theta)=\underline{u}+\frac{\tilde{u}+1}{2}\,\left(\bar{u}-\underline{u}\right),$ (15)
which maps $\tilde{u}\in[-1,\,1]$ onto $\underline{u},\,\bar{u}]$. Together, the construction 14 and de-normalization 15 ensures physical control input will always be $u^{*}$ at $x^{*}$ for any $\theta$ under the desired input constraints. The effectiveness of this simple strategy for enforcing input constraint is validated in Section IV.
III-B The Neural-NPV algorithm
Within Neural-NPV, we jointly train the Lyapunov network $\phi_{\textup{NN}}$ and controller network $\pi_{\textup{NN}}$ using a two stage procedure. In Stage I, we use counterexample guided synthesis to obtain a working controller and a Lyapunov function candidate over the full domain $\mathcal{X}\times\Theta$, where $\mathcal{X}$ is the region of interest. In Stage II, we refine the controller and Lyapunov networks using a level-set-guided loss to certify a robust ROA and a surrogate loss that promotes the growth of the robust ROA. The algorithm is summarized in 34.
III-B1 Stage I: Counterexample-guided joint synthesis
Existing work on learning certified control typically first generates samples from the state space, use these samples to evaluate the violation of the Lyapunov functions and optimizes the network parameters to minimize the violation. We follow this idea. However, here, we sample from the space of $x$, and $\theta$, and $\dot{\theta}$ i.e., $\mathcal{X}$, $\Theta$, and $\Omega$. At each iteration, we generate $N$ number of $x$-$\theta$ pairs, and then use CEGIS (Counter-example guided Inductive Synthesis) [1] framework to iteratively generate adversarial samples $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ using projected gradient descent (PGD) [23] to generate the worst adversarial examples that violate the following objective:
$\hskip-11.38109ptL_{\dot{V}}=\sum_{\dot{\theta}\in\textup{Ver}(\Omega)}\!\dot{V}(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}},\dot{\theta})+k_{3}\|x^{i}_{\textup{adv}}\|^{2}.$ (16)
By maximizing 16, we aim to find adversarial samples that have the maximum violation of the condition 9c. Since 9a and 9b has already been satisfied by construction as explained in Section III-A, we only need to evaluate 9c. At each $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$, we evaluate $V(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and $u(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and thus assess whether the conditions in 9 are violated at $(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})$ and if so, the level of violation. These adversarial examples are added to a buffer and then used to optimize the following empirical Lyapunov loss:
$\hskip-11.38109ptL_{lya}=\frac{1}{N}\sum_{i=1}^{N}\textup{ReLU}\!\!\!\!\!\!\sum_{\dot{\theta}\in\textup{Ver}(\Omega)}\!\!\!\!\left(\dot{V}(x^{i}_{\textup{adv}},\theta^{i}_{adv},\dot{\theta})+\!k_{3}\|x^{i}_{\textup{adv}}\|^{2}\right),$ (17)
which characterizes the violation of 9c using the $N$ samples. We minimize the Lyapunov loss in 17 to jointly learn a working controller and a Lyapunov function candidate over the ROI $\mathcal{X}\times\Theta$. We stop once we see the counterexample violation rate drops below $1\%$ and the counterexample buffer size plateaus for 100 consecutive iterations, indicating that the network satisfies condition 9c over the vast majority of the sampled state-disturbance space.
III-B2 Stage II: Level set-guided ROA refinement
In Stage I, we aim to get a working controller and a Lyapunov function candidate for the whole ROI. In Stage II, we introduce a level set
$\Lambda^{\rho}_{V}=\left\{x\in\mathbb{R}^{n}\mathrel{\mathop{\ordinarycolon}}\max_{\theta\in\Theta}\,V(x,\theta)\leq\rho\right\},$ (18)
in the training process, and refine the Lyapunov function candidate and the controller together, so that the level set represents a robust ROA, and furthermore its size is maximized. In practice, we can fix the constant $\rho$, e.g., to 1, since we can always update $V(x,\theta)$ to adjust the level set.
For $\Lambda^{\rho}_{V}$ to be a valid robust ROA, the Lyapunov conditions defined in 9 must be satisfied for all state inside $\Lambda^{\rho}_{V}$ and for all $\theta\in\Theta$. Given our parameterization, we only need to focus on 9c. Thus, we define the following loss function:
$\hskip-5.69054ptL_{\dot{V};\rho}\!=\frac{1}{N}\sum_{i=1}^{N}\!\min\left(L_{\dot{V}}(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}},\dot{\theta}),\rho\!-\!\max_{\theta\in\Theta}V(x^{i}_{\textup{adv}},\theta^{i}_{\textup{adv}})\right).$ (19)
We still follow the counterexample-guided training process and find adversarial examples in the level-set $\Lambda^{\rho}_{V}$ that violates 19. By concentrating at the level-set, we are able to significantly reduce the number of counterexamples and are able to find a robust ROA for a given NPV system. Given level-set value $\rho$ is fixed during the training procedure, the robust ROA found can be very conservative. In order to find a robust ROA that covers a large portion of the state space, we introduce a surrogate loss:
$\vskip-19.91692ptL_{\Lambda^{\rho}_{V}}=\frac{1}{N}\sum_{i=1}^{N}\max_{\theta\in\Theta}V(x^{i}_{\textup{{candidate}}},\theta^{i}_{\textup{{candidate}}})-\rho.$ (20)
Equation 20 indirectly promotes the growth of the level-set $\Lambda^{\rho}_{V}$. We sample uniformly from the state space to get ($x_{\textup{candidate}},\theta_{\textup{candidate}}$) that are used to compute 20. This surrogate loss ensures the robust level-set $\Lambda^{\rho}_{V}$ does not shrink to a neighborhood of the origin, and cover the region of interest as much as possible. The cost function of this refinement stage consists of both these objective functions. These two objectives play a complementary role, 19 certifies $\Lambda^{\rho}_{V}$ as a valid robust ROA while 20 tries to maximize its volume.
Algorithm 1 Neural-NPV Control
1:Lyapunov network $\phi_{\mathrm{NN}}$; Control network $\pi_{\mathrm{NN}}$ with initial parameters; Learning rate $\eta$; Region of interest $\mathcal{X}\times\Theta$; Counterexample (CEX) buffer $\mathcal{D}$; Candidate sample buffer $\mathcal{C}$; Maximum iteration $N_{\textup{iter}}$; Maximum epoch $N_{\textup{epoch}}$; PGD step $N_{\textup{PGD}}$; Max stagnation $n_{s}$; Samples per iteration $S$; Constant $k_{3}$; PGD step size $\beta$; Loss weights $\beta_{1}$ & $\beta_{2}$; level-set constant $\rho$
2:Stage I: Counterexample-guided joint synthesis
3:Initialize $\mathcal{D}$ $\leftarrow\varnothing$
[…]
5:  Sample randomly $S$ points ($x_{\textup{adv}},\theta_{\textup{adv}})\in(\mathcal{X}\times\Theta$)
6:  for $i=1\mathrel{\mathop{\ordinarycolon}}N_{\textup{PGD}}$ do
7:   Maximize violation via gradient ascent:
8:   ($x_{\textup{adv}},\theta_{\textup{adv}})\leftarrow\text{Project}_{(\mathcal{X}\times\Theta)}\Bigl((x_{\textup{adv}},\theta_{\textup{adv}})$ $\beta\nabla_{x_{\textup{adv}},\theta_{\textup{adv}}}L_{\dot{V}}(x_{\textup{adv}},\theta_{\textup{adv}},\dot{\theta},k_{3})\Bigr)$
9:  for every sample ($x_{\textup{adv}},\theta_{\textup{adv}})$ do:
10:   CEX validity check:
11:   if $L_{\dot{V}}(x_{\textup{adv}},\theta_{\textup{adv}})>0$ then:
[…]
14:   Compute $L_{lya}$ in 17 using all samples in $\mathcal{D}$
15:   $W_{V},W_{\pi}\;\leftarrow\;W_{V},W_{\pi}\;-\;\eta\,\nabla_{(W_{V},W_{\pi})}{L}_{lya}$
16:Stage II: level-set-guided ROA refinement
17:Initialize $\mathcal{D}$ $\leftarrow\varnothing$
[…]
19:  Sample randomly $S$ points ($x_{\textup{adv}},\theta_{\textup{adv}})\in(\mathcal{X}\times\Theta$)
[…]
21:   Maximize violation via gradient ascent in the level-set:
22:   ($x_{\textup{adv}},\theta_{\textup{adv}})\leftarrow\text{Project}_{(\mathcal{X}\times\Theta)}\Bigl((x_{\textup{adv}},\theta_{\textup{adv}})$ $\beta\nabla_{x_{\textup{adv}},\theta_{\textup{adv}}}L_{\dot{V};\rho}(x_{\textup{adv}},\theta_{\textup{adv}},\dot{\theta},k_{3};\rho)\Bigr)$
23:  for every sample ($x_{\textup{adv}},\theta_{\textup{adv}})$ do:
24:   CEX validity check:
25:   if $L_{\dot{V};\rho}(x_{\textup{adv}},\theta_{\textup{adv}})>0$ then:
[…]
27:  Sample randomly $S$ points ($x_{\textup{candidate}},\theta_{\textup{candidate}})\in(\mathcal{X}\times\Theta$), and append to buffer $\mathcal{C}$ $\leftarrow$ ($x_{\textup{candidate}},\theta_{\textup{candidate}})$
[…]
29:   Compute $L_{\dot{V};\rho}$ in 19 using all samples in $\mathcal{D}$
30:   Compute $L_{\Lambda^{\rho}_{V}}$ in 20 using all samples in $\mathcal{C}$
31:   $W_{V},W_{\pi}\leftarrow{W_{V},W_{\pi}}$ $\eta\nabla_{(W_{V},W_{\pi})}\Bigl(\beta_{1}\textup{ReLU}(L_{\dot{V};\rho})+\beta_{2}\textup{ReLU}(L_{\Lambda^{\rho}_{V}})\Bigr)$
32:  if CEX violation rate << 1% and no new CEX added to $\mathcal{D}$ for $n_{s}$ iterations then
33:   BREAK
34:Output $\phi_{\textup{NN}}$ & $\pi_{\textup{NN}}$
III-B3 Initialization of the control network
Before entering Stage I, we initialize our PD controller to enhance the training process. This is to avoid convergence issues when starting with a randomly initialized controller network. Inspired by [3], we first design a gain-scheduled linear state-feedback law based on LQR using local linearization at various $\theta$. We then train the initial controller network to match the gain-scheduled controller, i.e., we minimize the loss $\pi_{\textup{NN}}(x,\theta)-K(\theta)x||^{2}$ over $(x,\theta)$ samples so that the PD controller has a reasonable starting point.
III-C Empirical verification
In order to formally verify the Lyapunov conditions for our NPV systems, we would have to check not over only the augmented input $(x,\theta)\in(\mathcal{X}\times\Theta)$ but also for all cases $2^{n}_{\theta}\in\Omega$. The computational complexity involved in verifying Lyapunov conditions for NPV systems using formal verification tools such as $\alpha,\beta$-Crown very difficult. given our system limitations, we therefore adopt two empirical verification schemes to verify the Lyapunov conditions.
III-C1 PGD-based Verification
We run a strong adversarial PGD attack on our trained model, aiming to find states in the level-set $\Lambda^{\rho}_{V}$ that violates the Lyapunov derivative condition 9c (with a tolerance) for some $\theta\in\Theta$ and at any $\dot{\theta}\in\textup{ver}\Omega$. We generate $N_{\text{adv}}$ adversarial samples from this attack and consider our controller and Lyapunov function empirically verified when the violation rate, i.e., the ratio of samples that violate the condition 9c over all samples, remains below a small number, e.g., 1%, consistently.
III-C2 Trajectory-based Verification
We sample $10^{6}$ points from the level-set $\Lambda^{\rho}_{V}$ and see if the controller can converge all points inside $\Lambda^{\rho}_{V}$ under time varying $\theta(t)$. For each trajectory, we verify that the state converges to $x^{*}=0$ and that the Lyapunov function decreases monotonically, confirming that $\Lambda_{\rho}$ is a positively invariant set under the learned control law.
IV Simulation Results
We demonstrate the performance and scalability of our proposed methods through numerical experiments using two systems: a simple inverted pendulum and a quadrotor.
IV-A Inverted pendulum with varying control effectiveness
The dynamics of an inverted pendulum are given by
$\dot{x}(t)=\begin{bmatrix}\dot{\phi}(t)\\ \frac{g}{L}\sin(\phi(t))-\frac{b}{mL^{2}}\dot{\phi}(t)\end{bmatrix}+\theta(t)\begin{bmatrix}0\\ \frac{1}{mL^{2}}\end{bmatrix}u(t),$ (21)
where $x=[\phi,\dot{\phi}]^{T}$ and $u=\tau$ denote the state and control input (torque), respectively, $m=0.1$ $L=0.5$, $b=0.2$, and $g=9.81$ denote the mass, length, damping ratio, and gravitational constant. We impose an input constraint of $\mathinner{\!\left\lvert u(t)\right\rvert}\leq 3$ for all $t\geq 0$. The parameter $\theta(t)$ is introduced to characterize the control effectiveness of the actuator.
We assumed the following bounds: $\phi,\dot{\phi}]\in\mathcal{X}=[-\pi,\pi]\times[-6,6]$, $\theta(t)\in\Theta=[0.2,1]$ and $|\dot{\theta}(t)|\leq 0.1$. The structure of the Lyapunov and the controller network is shown in table Table I. Notice that the input to these networks are 4 channels rather than 3. In order to handle the discontinuity at $\phi=\pm\pi$ and provide a smooth, globally valid state representation for the network, we convert the angular state $\phi$ to $(\sin\phi,\,\cos\phi)$, which is fed to both control and Lyapunov networks together with $\dot{\phi}$ and $\theta$.
TABLE I: Network architectures and hyperparameters
Dynamical System $\phi_{\mathrm{NN}}$ dimension $\pi_{\mathrm{NN}}$ dimension $\eta$
Inverted Pendulum [4, 64, 128, 3] [4, 64, 128, 3] $10^{-5}$
3D Quadrotor [9, 64, 128, 6] [9, 64, 128, 64, 3] $10^{-5}$
We compared Neural-NPV to SOS-NPV [26], which jointly synthesizes a PD controller and PD Lyapunov function via SOS optimization, extended here with input constraints following [25]. From Fig. 2, we can see that Neural-NPV achieves a robust ROA whose volume is markedly larger than that obtained by SOS-NPV, and covers the whole state space along the $\phi$ axis. In SOS-NPV, the Lyapunov function was parameterized as $V(x,\theta)=x^{T}X^{-1}(\phi,\theta)x$, where $X(\phi,\theta)$ is a polynomial matrix function. Notice that we intentionally made $X(\phi,\theta)$ not dependent on $\dot{\phi}$ (whose derivative is directly affected by $u$) to have a convex optimization problem [14, 26]. In comparison, within Neural-NPV, the matrix term in the Lyapunov function 12 has a more general structure (not limited to the inverse of a polynomial function) and can depend on all states. This demonstrates the superiority of our framework compared to SOS-based optimization methods. To demonstrate the performance of the learned controller, we used a time-varying disturbance trajectory defined by $\theta(t)=0.6+0.4\cos(t)$. The resulting state trajectories and Lyapunov function evolutions are shown in Fig. 3 Trajectory-based verification confirmed $100\%$ convergence to the equilibrium across all $10^{6}$ sampled initial conditions. In our PGD-based verification experiments (with procedure detailed in Section III-C), we generated $10^{6}$ adversarial samples in $\Lambda^{\rho}_{V}\times\Theta$ and obtained a violation rate of approximately $1\%$ with a tolerance of $10^{-3}$, empirically validating $\Lambda^{\rho}_{V}$ as a robust ROA for the inverted pendulum system.
Figure 2: Robust ROA comparison between SOS-NPV and Neural-NPV with the same input constraint for the inverted pendulum system.
Figure 3: Trajectories of states from ten random initial points inside $\mathcal{X}$ under $\theta(t)=0.6+0.4\cos(t)$ (top), the Lyapunov function (middle), control input (bottom). Note that the input constraint $|u|\leq 3$ is respected for all cases.
Figure 4: 3D slice of the robust ROA $\Lambda^{\rho}_{V}$ for the quadrotor in $x$-$y$-$z$ space
Figure 5: Ten random trajectories of the state (top) and the corresponding Lyapunov function (bottom) for the quadrotor
Figure 6: The disturbance and the control input trajectories of the quadrotor. All the control inputs respect the input constraints (denoted by black dotted lines) explained in Section IV-B.
IV-B Quadrotor with varying external disturbance
To demonstrate the scalability of Neural-NPV to high-dimensional systems, we apply it to a quadrotor system subject to external disturbances:
[…]
where the state vector is $x=\begin{bmatrix}p_{x},\ p_{y},\ p_{z},\ \dot{p}_{x},\ \dot{p}_{y},\ \dot{p}_{z}\end{bmatrix}^{\top}\in\mathbb{R}^{6}$, and the scheduling parameter vector $\theta(t)=\begin{bmatrix}\theta_{x},\theta_{y},\theta_{z}\end{bmatrix}\in\mathbb{R}^{3}$ denotes the external disturbance forces along $x$, $y$, and $z$ axes. Here, we assume that the disturbance is known. In practice, the disturbances can be estimated using an adaptive estimator [22] or a disturbance estimator [4]. The attitude (roll, pitch, yaw) is parameterized as $(\phi,\theta_{p},\psi)$, and $\tau>0$ denotes the total thrust generated by the four rotors. The control input is defined as ${u}\triangleq\begin{bmatrix}\tau,\ \phi,\ \theta_{p}\end{bmatrix}^{\top}$. The outer-loop controller outputs desired thrust $\tau$ and attitude angles $(\phi,\theta_{p})$, which serve as references for an inner-loop attitude controller assumed to operate on a much faster timescale. As yaw is not actively controlled, we set $\psi=0$.
In 22, $g=9.81m/s^{2}$, $\mathbf{e}_{3}=[0,0,1]^{\top}$, and $R$ is the rotation matrix from body frame to the inertial frame following a ZYX (yaw-pitch-roll) sequence. The following state and input constraints are imposed: $(\phi,\,\theta_{p})\in[-90^{\circ},\,90^{\circ}]\times[-180^{\circ},\,180^{\circ}],\tau\in[0\,,\,11.3]\text{N}$. From 22, we can see that the dynamics are not affine in control. This prevents the direct application of SOS-based methods which are only applicable to control-affine systems [8, 10, 26].
We consider the following ROIs: $x\in\mathcal{X}=[-6,6]^{6},\theta(t)\in\Theta=[-2,2]^{3},\ \dot{\theta}(t)\in\Omega=[-0.5,0.5]^{3}$. The details of the Lyapunov network and the controller network can be found in Table I. To test the performance of the controller, we used a time-varying disturbance trajectory defined by $\theta(t)=\begin{bmatrix}2\sin(\pi t),2\cos(\frac{3}{4}\pi t),2\cos(\frac{\pi}{2}t)\end{bmatrix}^{T}$, which simulates an external disturbance acting on the quadrotor in the inertial frame. The resulting state trajectories and Lyapunov function evolution are shown in Fig. 5. From Fig. 6, we can see that after the states have already converged to the equilibrium in Fig. 5, the controller still constantly adjusts the control inputs to compensate for the external disturbances in order to keep the system at the equlibrium.
To verify our robust ROA empirically via the procedure described in Section III-C, we sampled $10^{6}$ trajectories in $\Lambda^{\rho}_{V}\times\Theta$. The controller drove the system to the equilibrium under all cases, empirically demonstrating the robustness of the controller. For PGD-based verification, we generated $10^{6}$ adversarial samples in $\Lambda^{\rho}_{V}\times\Theta$ and find that the counterexample violation rate obtained is approximately $1\%$ across all evaluation rounds with a tolerance of $10^{-3}$. Both of our empirical verification schemes validate $\Lambda^{\rho}_{V}$ as a robust ROA for the quadrotor system.
TABLE II: Empirical verification results under two evaluation schemes described in Section III-C
System Scheme Samples Violation rate Convergence
Inv. Pendulum PGD $0^{6} 1~1\%$
Trajectory $10^{6}$ 100%
Quadrotor PGD $0^{6} 11\%$
Trajectory $10^{6}$ 100%
V Conclusions
We presented Neural-NPV, a novel learning-based framework to jointly synthesize and verify a parameter-dependent (PD) Lyapunov function and PD controller for nonlinear parameter-varying (NPV) systems. Our two-stage training framework uses a counterexample-guided training procedure, and promotes the growth of the certified robust region of attraction. The framework is applicable to control-non-affine systems, is scalable to high-dimensional systems, and handles input constraints seamlessly. Numerical experiments on an inverted pendulum and a quadrotor system demonstrate the effectiveness of our proposed approach. One limitation of our current approach is the lack of formal verification of the learned systems. For future work, we intend to formally verify the learned controller and Lyapunov functions using neural network verifiers and experimentally validate the proposed approach on a real quadrotor system.